OpenStudy (anonymous):

Bill wanted to buy a slush at the store, so he checked his piggy bank. He only had dimes and nickels in it, and he had a total of 35 coins. He also had 5 more nickels than dimes. Which of the following slush sizes is the largest slush he could buy with his money? Show work. either a. Big- $1.89 b. Bigger - $2.49 c. Biggest - 2.79 d. Biggerest - $2.99

OpenStudy (anonymous):

b. Bigger - $2.49

OpenStudy (anonymous):

why is it b?

OpenStudy (anonymous):

let n = nickels and d = dimes n + d = 35 n - 5 = d n + (n -5) = 35 2n -5 = 35 2n = 40 n = 20 d = 15 .05 * n = 1.00 .10 * 15 = 1.50 1.00 + 1.50 = 2.50